xXMarcelieXx:

how do you find the range for this function? is there a way you can do it without graphing it ?

mhchen:

You can find the minimum and maximum of it by taking the derivative and setting where the derivative is 0.

xXMarcelieXx:

1 attachment
xXMarcelieXx:

part a

xXMarcelieXx:

can you show me how to do it

mhchen:

Well, I don't see where the question is.

xXMarcelieXx:

i posted above

imqwerty:

do you know what a square root is?

myeyeshurt:

The range of the function can be considered after determining the domain as this function is monotonic

myeyeshurt:

The domain is [3, infinity) and so the corresponding range is [k(3), k(infinity))=[0, infinity)

mhchen:

wow I'm blind, didn't even see that image

xXMarcelieXx:

how would you write the range in this form?

1 attachment
mhchen:

\(range(f) = \left\{ y \in [0,\infty) | y=f(x) \forall x \in [3,\infty)\right\}\) something like this I guess

xXMarcelieXx:

how would you do part a -c for the range ?

mhchen:

So we just did part (a) Part (b) is ln(2x-1) Its graph looks like this:  Doesn't the range look like all real numbers? Any number can be written as ln(2x-1) Here's a graph of sin(x):  It goes from -1 to 1 right? Now 3sin(x) would make it bigger 3 times.